OpenStudy (anonymous):

Name the property of real numbers illustrated by the equation. -4(x+3)=-4x-12 A. Commutative Property of Addition B. Associative Property of Addition C. Distributive Property D. Associative Property of Multiplication

OpenStudy (anonymous):

is this for homework?

OpenStudy (anonymous):

or a quiz/test?

OpenStudy (anonymous):

homework. its algebra two. and i just dont understand anything about it

OpenStudy (anonymous):

maybe it was because emily had nothing better to do on friday night than make up a question to post here

OpenStudy (anonymous):

WAITT, are you sure this is algebra 2. This seems pretty basic.

OpenStudy (anonymous):

@EmilyJernigan when you remove parentheses by multiplication, you use the distributive law (property)

OpenStudy (anonymous):

^^is true but the associative property also involved parentheses :/

OpenStudy (anonymous):

Positive! I can do almost anything but math!

OpenStudy (anonymous):

this says \(a(b+c)=ab+ac\) so in your example you have \[-4(x+3)=-4\times x+(-4)\times 3=-4x-12\]

OpenStudy (anonymous):

why do you explain so much....

OpenStudy (anonymous):

have nothing better to do on a friday night bwah xd

OpenStudy (anonymous):

ohhh! got it. :)

OpenStudy (anonymous):

good btw the distributive law is very important to remember because it links multiplication and addition it is also the one that is easiest to forget when doing other math problems

OpenStudy (anonymous):

Its been a while since ive done any math. And all my classes are online now. And i dont have any one to teach me. i gotta self teach. but all i have left is this class! but thank you. i got a couple more on this assignment

OpenStudy (anonymous):

go ahead and post. i promise not to ask "is this a homework"

OpenStudy (anonymous):

Solve the inequality. Graph the solution. 2|x+1/3|<9

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

this is going to take some work with fractions first divide both sides by 2 to get \[|x+\frac{1}{3}|<\frac{9}{2}\]

OpenStudy (anonymous):

then remove the parentheses and make a compound inequality \[-\frac{9}{2}<x+\frac{1}{3}<\frac{9}{2}\]

OpenStudy (anonymous):

then subtract \(\frac{1}{3}\) straight across to get \[-\frac{29}{6}<x<\frac{25}{6}\]

OpenStudy (anonymous):

okay. then i simplify?

OpenStudy (anonymous):

well there is nothing to simplify neither of the fractions reduce

OpenStudy (anonymous):

none of my answers look like that....? :/

OpenStudy (anonymous):

are you sure you copied the question down exactly right?

OpenStudy (anonymous):

 they look something like this.. this is answer A